How many numbers less than 300 can be formed from the digits 1,2,3,4 and 5
a) If repetitions are not allowed?
b) repetitions are allowed?

Answer by stanbon(57356)

How many numbers less than 300 can be formed from the digits 1,2,3,4 and 5
a) If repetitions are not allowed?
left-most digit 2 ways
Remaining 2 digits: 4C2 = 6
Total: 2*6 = 12 numbers
--------------------------------
b) repetitions are allowed?
left-most digit: 2 ways
Remaining 2 digits: 4^2 = 16
Total: 2*16 = 32 numbers
================================
Cheers,
Stan H.